# The properties of GSN 069 accretion disk from a joint X-ray and UV spectral analysis: stress-testing quasi-periodic eruption models

### v1

@@ 5 Conclusions
  • Using profile decomposition, we demonstrate that a combination of a point-source and diffuse stellar component well describes the innermost region of GSN 069’s FUV and optical images. The point source dominates the inner ($0.5\arcsec\times 0.5\arcsec$) region’s FUV flux, with the diffuse stellar contribution accounting for less than 5%. The diffuse component’s color is consistent with a moderately young stellar population (100–300 Myr). The high FUV intrinsic luminosity ($L_{\rm FUV}\gtrsim 10^{42}$ erg s-1) of the point-source precludes attributing any significant fraction of its flux to a nuclear stellar cluster (NSC), as this would require unphysically high star-formation surface densities.
@@
  • Analysis of the two epochs of UV spectra (1150–3150 Å) reveals that, after correcting for Galactic and intrinsic extinction/attenuation, the spectra are well described, to leading order, by $\nu L_{\nu}\propto\nu^{4/3}$, characteristic of the mid-frequency range of a standard thin disk. We also find that the integrated FUV flux decreased by $\sim$10from 2014 to 2018.
@@
  • We self-consistently model the X-ray and UV spectra together with the photometry of the diffuse stellar population. Our results demonstrate that a relatively simple disk model, with three free parameters, can simultaneously account for the X-ray emission and most (aside from the minor stellar contribution) of the UV spectra.
@@
  • Using a more realistic disk model, which incorporates relativistic corrections via numerical ray tracing, we constrain the disk inclination ($i$) to $31^{\circ}\lesssim i\lesssim 63^{\circ}$ (99% posterior). The black hole spin ($a$) remains ambiguous, with two equally likely modes in the posterior: (i) a highly spinning black hole ($a=0.95^{+0.02}_{-0.05}$) with $M_{\rm BH}=5.5^{+1.1}_{-0.6}\times 10^{6}~{}M_{\odot}$ and $i=59^{+4}_{-2}$ deg; or (ii) a slowly spinning black hole ($a=-0.19^{+0.48}_{-0.40}$) with $M_{\rm BH}=1.3^{+0.5}_{-0.3}\times 10^{6}~{}M_{\odot}$ and $i=41^{+7}_{-6}$ deg.
@@
  • In 2014, the outer disk radius ($R_{\rm out}$) was $\mathcal{O}(10^{3}R_{\rm g})$ for both solutions, a plausible size for a TDE disk if observed many years after disruption, though a fully time-dependent modeling of the source is still warranted. However, the evidence for simultaneous cooling ($T_{\rm p}^{2018}<T_{\rm p}^{2014}$) and expansion ($R_{\rm out}^{2018}>R_{\rm out}^{2014}$) of the disk are overwhelming. These are (unique) characteristics of a flow that is fed close to the black hole and expands viscously with a decreasing total inner disk mass; therefore, strong evidence for a TDE origin for the disk.
@@
  • Using the inferred system properties from both epochs ($M_{\rm BH}$, $\lambda_{\rm Edd}$, and $R_{\rm out}$), we test QPE model predictions. We argue that no published disk instability model can simultaneously produce an unstable disk (with the observed recurrence time) in 2018 while making the disk stable (no QPEs) in 2014, given $\lambda_{\rm Edd}^{2018}/\lambda_{\rm Edd}^{2014}$ is only $\approx 0.7$.
@@
  • For the disk/orbiter interactions models, although the large enough disk in 2018 is supportive of orbit disk collisions, the lack of the QPEs in 2014 when $R_{\rm out}$ was already $\gg R_{\rm orb}$ ($=30-100R_{\rm g}$), is hard to be reconciled with the model, unless a very fine-tune $\Sigma_{\rm disk}$ evolution has occurred. This possibility can be rigorously tested in future work via fully time-dependent modeling of the flow.
@@
  • In general, our work highlights that when multi-epoch high-quality multi-wavelength data is available and is physically and self-consistently analyzed, such analysis allows for rigorous testing of QPE models, which currently fail to reproduce the observables.

### v2

@@ 5 Conclusions
  • Using profile decomposition, we demonstrate that a combination of a point-source and diffuse stellar component well describes the innermost region of GSN 069’s FUV and optical images (see Fig. 1 and 2). The point source dominates the inner ($0.5\arcsec\times 0.5\arcsec$) region’s FUV flux, with the diffuse stellar contribution accounting for less than 5%. The diffuse component’s color is consistent with a moderately young stellar population (100–300 Myr). The high FUV intrinsic luminosity ($L_{\rm FUV}\gtrsim 10^{42}$ erg s-1) of the point-source precludes attributing any significant fraction of its flux to a nuclear stellar cluster (NSC), as this would require unphysically high star-formation surface densities (see §2.1).
@@
  • Analysis of the two epochs of UV spectra (1150–3150 Å) reveals that, after correcting for Galactic and intrinsic extinction/attenuation, the spectra are well described, to leading order, by $\nu L_{\nu}\propto\nu^{4/3}$, characteristic of the mid-frequency range of a standard thin disk. We also find that the integrated FUV flux decreased by $\sim$10from 2014 to 2018 (see Fig. 3).
@@
  • We self-consistently model the X-ray and UV spectra together with the photometry of the diffuse stellar population. Our results demonstrate that a relatively simple disk model, with three free parameters, can simultaneously account for the X-ray emission and most (aside from the minor stellar contribution) of the UV spectra (see Fig. 4).
@@
  • Using a more realistic disk model, which incorporates relativistic corrections via numerical ray tracing, we constrain the disk inclination ($i$) to $31^{\circ}\lesssim i\lesssim 63^{\circ}$ (99% posterior). The black hole spin ($a$) remains ambiguous, with two equally likely modes in the posterior (thought not value of the parameter space can be excluded): (i) a highly spinning black hole ($a=0.95^{+0.02}_{-0.05}$) with $M_{\rm BH}=7.5^{+1.0}_{-1.5}\times 10^{6}~{}M_{\odot}$ and $i=59^{+4}_{-2}$ deg; or (ii) a slowly spinning black hole ($a=-0.19^{+0.48}_{-0.40}$) with $M_{\rm BH}=1.5^{+0.5}_{-0.3}\times 10^{6}~{}M_{\odot}$ and $i=41^{+7}_{-6}$ deg (see Fig. 5 and 6). These bimodal posterior distribution, maybe be broken, by e.g., adding more X-ray spectra into the fitting.
@@
  • In 2014, the outer disk radius ($R_{\rm out}$) was $\mathcal{O}(10^{3}R_{\rm g})$ for both solutions, a plausible size for a TDE disk if observed many years after disruption, though a fully time-dependent modeling of the source is still warranted. However, the evidence for simultaneous cooling ($T_{\rm p}^{2018}<T_{\rm p}^{2014}$) and expansion ($R_{\rm out}^{2018}>R_{\rm out}^{2014}$) of the disk are overwhelming (see Fig. 6, 7). This behavior points to the fact that GSN 069’s disk is not in a steady-state, and inferred evolution is (unique) characteristic of a flow that is fed close to the black hole and expands viscously with a decreasing total inner disk mass; therefore, strong evidence for a TDE origin for the disk (see 4.1).
@@
  • Using the inferred system properties from both epochs ($M_{\rm BH}$, $\lambda_{\rm Edd}$, and $R_{\rm out}$), we test QPE model predictions. We argue that no published disk instability model can simultaneously produce an unstable disk (with the observed recurrence time) in 2018 while making the disk stable (no QPEs) in 2014, given $\lambda_{\rm Edd}^{2018}/\lambda_{\rm Edd}^{2014}$ is only $\approx 0.7$ (see §4.2.1).
@@
  • For the disk/orbiter interactions models, although the large enough disk in 2018 is supportive of orbit disk collisions, the lack of the QPEs in 2014 when $R_{\rm out}$ was already $\gg R_{\rm orb}$ ($=30-200R_{\rm g}$), is hard to be reconciled with the model, unless a very fine-tune $\Sigma_{\rm disk}$ evolution has occurred (see §4.2.2). This possibility can be rigorously tested in future work via fully time-dependent modeling of the flow.
@@
  • In general, our work highlights that when multi-epoch high-quality multi-wavelength data is available and is physically and self-consistently analyzed, such analysis allows for rigorous testing of QPE models, which currently struggle to reproduce all the observables simultaneously.